# Can GenAI Move from Individual Use to Collaborative Work? Experiences, Challenges, and Opportunities of Integrating GenAI into Collaborative Newsroom Routines

### v2

@@ 4.3. Data Analysis
We analyzed the data using reflexive thematic analysis (Braun and Clarke, 2019). This approach enabled us to identify recurring patterns across the dataset while remaining attentive to the contextual specificities of participants’ lived experiences. Consistent with many practices in qualitative HCI research (Chinh et al., 2019; Xiao et al., 2025b), we did not conduct statistical measures of inter-coder reliability. Our methodological stance is grounded in an interpretivist paradigm, which views knowledge as context-dependent and socially constructed. The goal was not to generate statistically generalizable results but to develop situated understandings that reflect participants’ perspectives (Morgan and Nica, 2020). All coding and theme development was conducted collaboratively by three researchers. Continuous discussions served to refine codes and themes, enrich interpretations, and ensure that our conclusions were jointly reasoned and methodologically sound. In the following section 5, section 6, and section 7, we outline the key themes that emerged from our analysis. These findings themes are also summarized in Figure 1 for reference.
@@
5. Findings: Still Individual Use, but with Expectations for Integrating GenAI into Collaboration (RQ1)
@@
Our findings reveal that although GenAI use is already widespread at the individual level, it is highly personalized, informal, and largely invisible to colleagues (subsection 5.1). Yet, these individualized practices have not yet coalesced into team-based routines or newsroom-wide standards for future collaborations. Even so, many journalists described actively working to ensure that their own GenAI use, and that of their teammates, remained aligned with professional values in collaborative work by individual efforts. Despite the lack of formal coordination, journalists nonetheless express a strong desire for GenAI to become a part of their collaborative routines (subsection 5.2).
5.1. Widespread but Individualized Use of GenAI
@@
Across our interviews, participants expressed a wide spectrum of attitudes toward the use of GenAI in the newsroom, ranging from pessimistic to optimistic. Some viewed GenAI as a potential threat to journalistic rigor, fearing that it might “destroy the entire editorial process, make journalists lazier and less reflective, and ultimately lead them to rely on GenAI-generated text at the expense of their own writing skills” (P22). Others held a more hopeful outlook, believing that “GenAI can significantly improve journalistic productivity by saving time on tedious tasks and freeing up energy for meaningful investigative work” (P25).
Despite these differences, there was broad consensus on one point: the use of GenAI by individual journalists is already widespread and likely irreversible. “It’s nearly impossible to stop individual journalists from using GenAI,” P15 noted. “The best we can do is figure out how to manage it in the organizational level.” P15 further contextualized this trend by drawing an analogy to broader societal shifts. “It’s not just journalists, everyone is using some form of GenAI nowadays,” P15 explained. “Even eight-year-old child today is using GenAI. If school teachers cannot stop the child from using it, how could we possibly expect to stop a professional adult journalist?”
@@
P2, a senior manager in her newsroom who was responsible for major organizational decisions and frequent communication with government authorities, cited a 2024 report by the research team from Xinhua News Agency (Xinhua Research Institute, 2024), the China’s most prominent state newspaper, stating that “very few newsrooms have expressed total rejection of GenAI in organizational level. Most are simply waiting for the right moment, for the GenAI technology to mature enough to be formally introduced into collaborative newswork.”
In P2’s view, whether GenAI should eventually be formally integrated into newsroom routines is hardly a matter for debate, since the “widespread individualized use of GenAI” by journalists has already made its presence in newswork effectively irreversible; the real question is how this integration ought to take place. She emphasized that the issue was not contingent on any individual journalist’s pessimism or outright rejection of GenAI, nor was it realistically possible for newsrooms either to prohibit individual journalists from using GenAI or to reject the technology entirely at the organizational level. After all, P2 believed, individual journalists could always access GenAI on their personal phones or computers without informing the organization. As long as a subset of journalists within the organization were already making use of GenAI in their everyday work, skepticism at either the personal or organizational level could not meaningfully halt its diffusion.
@@
It is also worth noting that in our study, all direct interactions between journalists and GenAI took the form of conversational exchanges with chat-based GenAI systems such as Qwen or DeepSeek. Participants typically engaged GenAI by typing prompts in natural language, requesting draft texts, summaries, translations, or stylistic adjustments. In other words, GenAI was used as an interactive writing partner rather than through more technical means such as API calls, system integration, or backend customization. This use of GenAI illustrates its deep overlap with the core tasks of news production, integrating directly into the writing and editing, while also reinforcing its status as a largely private, individualized practice that can be easily concealed from colleagues and editors.
5.2. Motivations for GenAI Integrations into Collaborative Newsroom Routines
@@
In our sampled newsrooms, maintaining the alignment between GenAI usage and journalistic expectations currently depends almost entirely on individual effort. In newsroom practice, front-line journalists typically draft stories, exchange leads with peers, and submit copy to editors, who then review, revise, and make publication decisions. Within this layered process, front-line journalists who experiment with GenAI must not only ensure the appropriateness of their own outputs but also anticipate how these outputs will intersect with editors’ judgments and colleagues’ contributions. Most participants described the extra effort this entailed: double-checking GenAI-generated text before passing it along, quietly monitoring whether teammates might have relied on GenAI, and compensating for possible errors to avoid undermining collective products. Editors, too, were not merely gatekeepers reacting to GenAI-produced content; some such as P5 and P7 also adopted GenAI tools to support their own editorial work, for example, using them to generate alternative headlines. These editorial uses similarly required individualized scrutiny to ensure alignment with house style and professional standards.
Yet such vigilance often left participants feeling isolated and burdened. Despite journalists’ increasing comfort using GenAI on their own terms, many participants expressed uncertainty about how their use intersected with that of their colleagues. Few had insight into how others were deploying GenAI in daily work, and most described their usage as “private” (e.g., P9, P13, P14) or “informal” (e.g., P5, P7, P16), a tool kept largely to themselves rather than openly discussed or coordinated within teams.
This dynamic not only reveals the professionalism and ethical awareness of individual journalists, but also underscores an urgent need for collective mechanisms that can support the responsible use of GenAI. Several participants, including P2, P4, P6, P8, P14, and P22, expressed deep concerns that the lack of formal integration of GenAI into newsroom workflows amplifies risks in collaborative work. Without shared guardrails or institutional guidance, they feared that GenAI could introduce errors into collective news products and causes team or even organizational level harms.
At the same time, participants saw promise in the idea that integrating GenAI into collaborative routines could help distribute responsibility more equitably across teams and reduce the cognitive and moral load currently borne by individual journalists. P12, for instance, explicitly questioned why no newsroom-wide policies or tools had been introduced, suggesting that management’s delay in institutionalizing GenAI might be a strategic way to offload risk onto individuals while avoiding organizational accountability.
@@
Many participants (e.g., P3, P12 and P20) expressed a strong desire for more open conversations about the future role of GenAI in collaborative newswork. Rather than leaving the GenAI tool in a gray zone of informal, backstage experimentation for individuals, journalists hoped to see it open to scrutiny. They envisioned shared discussions about best collaborative practices, institutional expectations, and ethical tradeoffs as necessary foundations for building trust, enabling peer learning, and clarifying the legitimate place of GenAI in teamwork.
Notably, participants frequently contrasted GenAI with earlier personal productivity tools such as Microsoft Word or Excel. While these tools are similarly used by individual journalists, they were seen as passive instruments for documentation or formatting, with “no agency in shaping the content itself” (P7). GenAI, by contrast, was understood to be an active agent, generating, suggesting, and transforming journalistic text. Unlike word processors that could be mostly individualized, GenAI must be understood, coordinated, and made transparent within the newsroom collective.
In short, moving GenAI from the private domain into the realm of collective routines emerged as a widely shared aspiration across the newsroom contexts we studied.
6. Findings: Challenges in Integrating GenAI into Collaborative Newsroom Routines (RQ2)
@@
In the section 5, we demonstrated that the current use of GenAI in newsrooms is highly individualized. Journalists engage with GenAI tools in personal, often private ways, with limited visibility into how their colleagues are using similar technologies. This fragmented usage stands in sharp contrast to the collaborative nature of journalistic work, which typically relies on shared routines, editorial review, and team-based production. In this section, we investigate why establishing collaborative GenAI routines has proven especially difficult in newsroom settings. It is important to clarify that our findings focus only on traditional newswork and do not extend to journalist-technical companies or journalists-AI engineers partnerships. This is because the technical teams or companies building GenAI tools do not participate in the day-to-day newswork.
@@
These challenges in integrating GenAI into collaborative newsroom routines can be broadly grouped into two interrelated domains: structural and cultural. Structurally, many newsrooms lack formal coordination mechanisms for GenAI usage, making it difficult to establish shared visibility or accountability across team members (subsection 6.1). Culturally, journalists often regard their GenAI use as private, experimental, or even stigmatized, leading to reluctance in disclosing their practices or outputs to peers and supervisors (subsection 6.2).
@@ 6.1. Structural Challenges: Lack of Organizational Mechanisms for Integration
While individual journalists have actively experimented with GenAI tools, efforts to embed these practices into newsroom-level collaborative routines face critical structural barriers. These structural challenges stem from a lack of organizational incentives or mandates to coordinate GenAI usage (subsubsection 6.1.1), insufficient technical infrastructure to support collaborative workflows (subsubsection 6.1.2), and heightened public scrutiny that discourages formal integration without robust safeguards (subsubsection 6.1.3).
@@
Although many newsroom leaders (e.g., P1, P2, P3) are aware that individual journalists within their organizations are already experimenting with GenAI tools, they rarely translate this awareness into concrete action at the team level. “We’ve done a lot to build internal GenAI systems, which is mainly to satisfy government expectations,” P13 explained, “but unfortunately, we don’t yet have a coherent collaboration guide about how to use GenAI for our journalists. No one really knows how it could enhance teamwork.” From a broader organizational perspective, P1, a senior manager, articulated this divide more explicitly. P1 emphasized that newsroom-level initiatives, often launched in response to national calls, tend to prioritize the development of specific GenAI systems, rather than paying close attention to how journalists incorporate GenAI into their daily practices.
What participants described as “GenAI systems” typically referred to newsroom-specific projects that fine-tuned existing large language models or built customized front-end interfaces to showcase AI-powered news functions. These prototypes, ranging from tools that automatically drafted short news items, to those generating video subtitles, summarizing interview transcripts, or retrieving archival content, functioned less as integrated aids to newsroom collaboration and more as deliverables that could be presented to authorities or internal management as evidence of technological advancement.
@@
This gap between symbolic system-building and actual newsroom routines reflects a deeper uncertainty about how GenAI should be integrated into the labor structures of journalism. Several editors (n=7) voiced explicit concerns that formalizing GenAI use might unintentionally signal an endorsement of automation, thereby threatening long-standing norms around authorship, creativity, and editorial authority. P13 expressed this dilemma: “If we say it’s okay to use GenAI, would that encourage journalists to rely on it too much, instead of their own judgment? And how would our audience perceive this? Maybe they’d think, ‘Oh, this newspaper is just letting AI write the news now,’ or ‘Their journalism must be so poor that AI can step in.’ Some news audiences still carry a lot of negative impressions about GenAI.” This institutional hesitation creates a structural vacuum: GenAI becomes a tolerated but unregulated practice, operating in a gray zone between tacit acceptance and formal prohibition.
@@
Moreover, even in organizations that recognize the potential value of collaborative GenAI use, many newsrooms lack the expertise or confidence to define what such collaboration should look like. As senior manager P3 put it bluntly: “We don’t even know how to try. What does GenAI collaboration actually mean?” P3 contrasted this ambiguity in collaboration with the current state of highly individualized GenAI use, where accountability remains clear. “Look, if everyone is using GenAI in a personal and unofficial way, and something goes wrong, we know who to hold responsible: it’s the individual journalist, right?” However, P3 raised concerns about what might happen if the newsroom were to formalize guidelines for how GenAI should be integrated into team workflows. “Now imagine we create an official framework for collaboration between human journalists and GenAI. Let’s say we even publish a clear policy about how GenAI shall be incorporated into teamwork. Then if a problem arises, who’s accountable? The journalist might say, ‘I was just following the policy.’ So where does that leave us?”
P3’s observation points to a critical tension: any move to institutionalize GenAI collaboration may inadvertently blur lines of responsibility, making it harder for editors or managers to attribute fault when mistakes occur. In contrast, the current hands-off approach, where GenAI use is quietly tolerated but not officially endorsed, preserves individual accountability and limits organizational liability. This ambiguity creates a paradoxical situation. On the one hand, organizations hesitate to create collaborative structures for fear of enabling blame deflection and institutional risk. On the other hand, without formal coordination mechanisms, GenAI use remains fragmented and inconsistent, inhibiting the development of scalable, transparent, and trustworthy practices. As a result, many newsrooms remain trapped in a structural limbo, aware of GenAI’s potential to reshape workflows, yet unwilling to define what responsible collaboration actually entails.
@@
6.1.2. Second, the technical infrastructure of most Chinese newsrooms poses significant challenges to building collaborative GenAI routines.
@@
Most GenAI tools were accessed individually via external platforms, such as Qwen Chat, Doubao, or DeepSeek. There are no built-in mechanisms in newsrooms to track when, where, or how GenAI has been used in the production of a specific story. Editors and collaborators often lack visibility into which portions of a draft were generated by AI, making it more difficult to assess the originality or editorial appropriateness of the content. As P2 explained, “People are mostly using GenAI on external platforms, which makes it hard for us to trace how they’re using it. We’re trying to build in-house GenAI systems with technical companies, but those are still in pilot stages. In most cases, journalists are still using external GenAI tools.”

### v3

@@ 4.3. Data Analysis
We analyzed the interviews using reflexive thematic analysis (Braun and Clarke, 2019), following an iterative, interpretive, and team-based approach. Our analytic process unfolded across three stages.
Stage 1: Immersion and initial coding. The researchers read the transcripts in full to build familiarity with the dataset. Three researchers independently conducted open coding on a shared subset of transcripts, noting instances of GenAI use, collaboration challenges, informal workarounds, value tensions, and moments of uncertainty or ambiguity. These codes were intentionally descriptive and numerous, capturing as much detail as possible without imposing early structure. The team then met to compare initial codes, discuss divergent interpretations, and collaboratively construct a preliminary coding frame that reflected our emerging analytic interests.
Stage 2: Systematic coding and iterative refinement. Using the initial coding frame, the full dataset was coded by the research team. Following the principles of reflexive qualitative analysis, disagreement was treated not as a reliability problem but as a productive site for interpretation (Morgan and Nica, 2020). Researchers wrote analytic memos to trace emerging patterns, theoretical questions, and surprising contradictions in participants’ accounts. We met regularly to refine code definitions, merge overlapping codes, elevate promising ones, and prune those that did not hold analytically. These discussions helped us critically examine our assumptions and ensure that our interpretations remained grounded in, rather than imposed onto, the data.
Stage 3: Theme construction and validation. Once coding was complete, we clustered related codes into broader themes that captured recurrent patterns across participants’ experiences. In line with reflexive thematic analysis, themes were understood as meaning-based patterns, not frequency counts. We repeatedly returned to raw excerpts to validate the coherence and distinctiveness of each theme, testing whether they adequately represented participants’ experiences and reflected the tensions between individual GenAI experimentation and collaborative newsroom workflows. The final themes, detailed in section 5, section 6, and section 7, are also summarized in Figure 1.
4.4. Positionality Statement
We acknowledge that our backgrounds, methodological training, and prior professional experiences shape how we approached this study and interpreted participants’ accounts. Our research team is trained in Human–Computer Interaction, Communication, Journalism, and Organizational Studies, and several team members have prior work experience in digital journalism and human–AI collaboration. Two researchers previously worked in or closely studied Chinese newsrooms, which gives us familiarity with the professional norms, political sensitivities, and hierarchical structures that characterize this environment.
Throughout the study, we engaged in deliberate reflexive practice, following approaches in HCI and CSCW that encourage researchers to critically examine their assumptions throughout the research process (Kuo et al., 2023). We continuously asked: (1) How might our own expectations about GenAI influence the questions we asked or the patterns we foregrounded? (2) What biases do we bring from our familiarity with newsroom culture or from our broader research agendas on human–AI collaboration? (3) Who benefits from this research, and whose perspectives might remain underrepresented? Each team member documented these reflections during analysis, and we compared them during coding meetings. These reflexive discussions made us more attentive to the multiple interpretations of journalists’ practices—for example, recognizing that cautious GenAI use could stem not only from technical distrust but also from hierarchical accountability, political sensitivity, or cultural norms around authorship. We aimed to conduct our research with humility, respect for participants’ expertise, and an awareness of how our positionality shaped the knowledge we produced.
5. Findings: Still Individual Use, but with Expectations for Coordinating GenAI into Collaboration (RQ1)
@@
Our findings reveal that although GenAI use is already widespread at the individual level in our sampled newsrooms, it is highly personalized, informal, and largely invisible to colleagues (subsection 5.1). Yet, these individualized practices have not yet coalesced into teamwork or newsroom-wide standards for future collaborations. Even so, many journalists described actively working to ensure that their own GenAI use, and that of their teammates, remained aligned with professional values in collaborative work by individual efforts. Despite the lack of formal coordination, journalists nonetheless express a strong desire for GenAI to become a part of their collaborative newswork (subsection 5.2).
5.1. Widespread but Individual Use of GenAI in Chinese Newsrooms
@@
Across our interviews, participants expressed a wide spectrum of attitudes toward the use of GenAI in the newsroom, ranging from pessimistic to optimistic. Some viewed GenAI as a potential threat to journalistic rigor, fearing that it might “destroy the entire editorial process, make journalists lazier and less reflective, and ultimately lead them to rely on GenAI-generated text at the expense of their own writing skills” ($P22_{\text{male, journalist}}$). Others held a more hopeful outlook, believing that “GenAI can significantly improve journalistic productivity by saving time on tedious tasks and freeing up energy for meaningful investigative work” ($P25_{\text{female, journalist}}$).
@@
Despite these differences, there was broad consensus on one point: the use of GenAI by individual journalists is already widespread in 2025 in China and likely irreversible. $P2_{\text{female, editor/manager}}$, a senior manager in her newsroom who was responsible for major organizational decisions and frequent communication with government authorities, cited a 2024 report by the research team from Xinhua News Agency (Xinhua Research Institute, 2024), the China’s most prominent state newspaper, stating that “very few newsrooms have expressed total rejection of GenAI in organizational level. Most are simply waiting for the right moment, for the GenAI technology to mature enough to be formally introduced into collaborative newswork.”
$P15_{\text{female, journalist}}$ offered a complementary perspective grounded in her everyday work. “As far as I know, many of my colleagues are already using GenAI. We use it to brainstorm and to revise our early drafts. For me, the final article is nearly no similarity compared to the AI draft because I revise it so much, but GenAI is essential in the very beginning, when getting the first version down is the hardest part for me.” In her view, when journalists talk about the ‘widespread’ use of GenAI, this does not mean that everyone uses GenAI all the time, nor that GenAI is involved in every story. Rather, “people inevitably find themselves turning to GenAI at certain moments,” as she described.
Journalists’ engagement with GenAI was selective, situational, and guided heavily by personal judgment. Some (e.g., $P19_{\text{male, journalist}}$, $P20_{\text{male, journalist}}$) relied on GenAI multiple times a day for drafting or summarizing; others used it only when pressed for time (e.g., $P17_{\text{male, journalist}}$), and a few avoided it entirely for tasks they viewed as requiring professional integrity or narrative control (e.g., $P21_{\text{male, journalist}}$). Among our participants, they commonly felt, most journalists draw on GenAI when they feel it is needed, yet such on-demand use is still widespread across individuals.
Yet it is precisely this fragmented and uneven individualized use that emerged as consequential in our study. In a highly interdependent newsroom environment, the core challenge was the pervasive uncertainty it created: journalists could rarely tell whether, when, or how their colleagues had used GenAI. This ambiguity set off what many participants described as a subtle chain of mutual suspicion. As $P7_{\text{male, editor}}$ put it, “Sometimes I try to edit a draft and I really wonder: did they use GenAI here or not? I have no way to know, but I have to check it as if they did.” Journalists were unsure whether drafts they received from peers contained undisclosed AI-generated passages; editors often wondered whether a stylistic shift or factual slip reflected GenAI involvement; and teammates speculated about how much additional verification they were expected to perform. In the subsection 5.2, we show how these uncertainties and suspicion chains motivated journalists to imagine ways of bringing individualized GenAI practices into more coordinated and transparent forms of collaborative and organizational coordination.
@@
It is worth noting that in our study, all direct interactions between journalists and GenAI took the form of conversational exchanges with chat-based GenAI systems such as Qwen or DeepSeek. Participants typically engaged GenAI by typing prompts in natural language, requesting draft texts, summaries, translations, or stylistic adjustments. In other words, GenAI was used as an interactive writing partner rather than through more technical means such as API calls, system integration, or backend customization. This use of GenAI was integrated directly into the writing and editing, while also reinforcing its status as a largely private, individualized practice that can be easily concealed from colleagues and editors.
It is also important to note that, despite the rapid expansion of multimodal GenAI capabilities, participants overwhelmingly described using GenAI primarily for text-based tasks rather than image or video generation. This pattern was not due to a lack of technical availability, as many journalists were familiar with GenAI image and video tools but reflected deeper professional and regulatory constraints. In Chinese newsrooms, AI-generated images or videos are largely considered inappropriate for routine news production and are only acceptable when intentionally staged as AIGC content or published within clearly experimental formats (e.g., special AI zones). Outside of these contexts, AI-generated visuals are widely viewed as carrying political, ethical, and credibility risks: they can be mistaken for manipulated media, trigger public suspicion, or violate strict editorial norms around visual authenticity. As several participants noted, “AI pictures are a taboo in everyday news, we are not allowed to use that in our work without permission” ($P11_{\text{female, journalist}}$) unless their artificial nature is the point of the story. By contrast, text generation was seen as easier to regulate, revise, and ‘humanize’ through subsequent editing, making it the dominant modality through which GenAI entered everyday newswork.
5.2. Expectation for Better Coordinating GenAI into Collaborative Newswork
@@
Across the newsrooms we studied, journalists firstly described a shared sense that GenAI could no longer purely remain an individualized, private tool because this arrangement placed disproportionate burdens on individual journalists and editors. In our sampled newsrooms, maintaining the alignment between GenAI usage and journalistic expectations currently depends almost entirely on individual effort. Generally, front-line journalists typically draft stories, exchange leads with peers, and submit copy to editors, who then review, revise, and make publication decisions. Within this layered process, journalists experimenting with GenAI could not only ensure the appropriateness of their own outputs but also anticipate how these outputs will intersect with editors’ judgments and colleagues’ contributions.
Most participants described the extra effort this entailed: double-checking GenAI-generated text before passing it along, quietly monitoring whether teammates might have relied on GenAI, and compensating for possible errors to avoid undermining collective products. Editors, too, were not merely gatekeepers reacting to GenAI-produced content; some such as $P5_{\text{female, editor}}$ and $P7_{\text{male, editor}}$ also adopted GenAI tools to support their own editorial work. In both roles, however, using GenAI created a dual layer of responsibility: journalists and editors had to remain accountable for the AI-assisted text they produced themselves, while also devoting additional effort to scrutinizing the materials they received from colleagues. According to $P7_{\text{male, editor}}$, “this scrutiny now extended to stylistic and tonal indicators, especially given the recent rise of public vigilance toward AI-generated writing”. $P7_{\text{male, editor}}$ noted, “even minor shifts in phrasing could draw audience suspicion, placing heightened pressure on editors, who hold final responsibility of any AI-related risks before publication.”
Participants further emphasized that GenAI’s largely private and disconnected use created pervasive uncertainty in collaborative work, reinforcing their desire for more coordinated practices. Few had insight into how others were deploying GenAI in daily work, and most described their usage as “private” (e.g., $P9_{\text{male, editor}}$, $P13_{\text{male, editor/journalist}}$, $P14_{\text{female, editor/journalist}}$) or “informal” (e.g., $P5_{\text{female, editor}}$, $P7_{\text{male, editor}}$, $P16_{\text{female, editor}}$), a backstage tool rather than something openly discussed within teams. This produced hidden coordination labor: journalists and editors worried about how much extra verification was needed, and whether colleagues had unknowingly relied on GenAI in ways that could introduce subtle errors.
Moreover, because these uncertainties threatened not just individual journalists but team- and organizational-level outcomes, participants argued that GenAI required collective guardrails. Several participants (n=11), described deep concerns that the lack of shared standards allowed GenAI to introduce errors that would impact collective news products and organizational credibility. Front-line journalists worried that undisclosed AI use could lead to “mistakes being traced back” to them, exposing them to “individual blame” and “disciplinary scrutiny” ($P14_{\text{female, editor/journalist}}$); editors feared reputational damage because they are “ultimately accountable for publication quality” ($P8_{\text{male, editor}}$); managers were concerned about political, regulatory, and ethical consequences if AI-assisted content contaminated organizational outputs, “especially in China that are so sensitive, even small mistakes can become dangerous” ($P1_{\text{female, editor/manager}}$).
These perceived risks also highlighted a broader issue of fairness, motivating participants to call for more explicit organizational practices that would distribute responsibility more equitably. Many felt that the current ambiguity enabled organizations to “offload risks onto individuals” ($P12_{\text{female, journalist}}$), especially junior journalists and editors, while avoiding institutional accountability.
@@
Finally, participants hoped that bringing GenAI into collective newswork would foster transparency, peer learning, and shared norms needed for responsible use. Many (e.g., $P3_{\text{male, editor/manager}}$, $P12_{\text{female, journalist}}$, $P20_{\text{male, journalist}}$) called for open conversations about GenAI’s future role in newswork, rather than allowing it to remain in a gray zone of private experimentation. They envisioned shared discussions about collaborative best practices, institutional expectations, and ethical tradeoffs as foundational for building trust and making GenAI a legitimate part of teamwork.
In short, participants’ motivations for coordinatingg GenAI into collaborative newswork stemmed from a combination of individual burdens, hidden coordination work, perceived organizational risks, concerns about fairness, and the need for shared norms.
6. Findings: Challenges in Coordinating GenAI into Collaborative Newswork (RQ2)
@@
In the section 5, we demonstrated that the current use of GenAI in newsrooms is highly individualized. Journalists engage with GenAI tools in personal, often private ways, with limited visibility into how their colleagues are using similar technologies. This fragmented usage stands in sharp contrast to the collaborative nature of journalistic work, which typically relies on shared norms, editorial review, and team-based production. In this section, we investigate why current Chinese newsrooms not yet developed mechanisms to coordinate. It is important to clarify that our findings focus only on traditional newswork and do not extend to journalist-technical companies or journalists-AI engineers partnerships.
@@
These challenges in coordinating GenAI into collaborative newswork can be broadly grouped into two interrelated domains: structural and cultural. Structurally, many newsrooms lack formal coordination mechanisms for GenAI usage, making it difficult to establish shared visibility or accountability across team members (subsection 6.1). Culturally, journalists often regard their GenAI use as private, experimental, or even stigmatized, leading to reluctance in disclosing their practices or outputs to peers and supervisors (subsection 6.2).
@@ 6.1. Structural Challenges: Lack of Organizational Mechanisms for Integration
While individual journalists have actively experimented with GenAI tools, efforts to embed these practices into newsroom-level collaborative norms face critical structural barriers. These structural challenges stem from a lack of organizational incentives or mandates to coordinate GenAI usage (subsubsection 6.1.1), insufficient technical infrastructure to support collaborative workflows (subsubsection 6.1.2), and heightened public scrutiny that discourages formal coordination without robust safeguards (subsubsection 6.1.3).
@@
In the context of China’s news industry, organizational priorities are often oriented toward meeting political expectations, demonstrating technological capability, and maintaining institutional legitimacy. Within this configuration, GenAI is primarily evaluated as a strategic or symbolic asset, used to signal innovation, efficiency, and alignment with policy agendas, rather than as a technology requiring sustained coordination within collaborative newswork. As a result, questions of how GenAI should be collectively governed, integrated into team-based workflows, or aligned with shared editorial norms have received comparatively little attention. Participants noted that this emphasis on institutional signaling over collaborative coordination leaves front-line teams to navigate GenAI-related risks individually, without clear, collectively articulated guardrails.
@@
Against this backdrop, although many newsroom leaders (e.g., $P1_{\text{female, editor/manager}}$, $P2_{\text{female, editor/manager}}$, $P3_{\text{male, editor/manager}}$) are aware that individual journalists within their organizations are already experimenting with GenAI tools, they rarely translate this awareness into concrete action at the team level. “We’ve done a lot to build internal GenAI systems, which is mainly to satisfy government expectations,” $P13_{\text{male, editor/journalist}}$ explained, “but unfortunately, we don’t yet have a coherent collaboration guide about how to use GenAI for our journalists. No one really knows how it could enhance teamwork.” From a broader organizational perspective, $P1_{\text{female, editor/manager}}$, a senior manager, articulated this divide more explicitly. $P1_{\text{female, editor/manager}}$ emphasized that newsroom-level initiatives, often launched in response to national calls, tend to prioritize the development of specific GenAI systems, rather than paying close attention to how journalists incorporate GenAI into their daily practices.
@@
This symbolic–practical gap also highlights a deeper structural hesitation: many organizations remain uncertain about what it would mean to meaningfully integrate GenAI into journalistic labor. Several editors (n=7) voiced explicit concerns that formalizing GenAI use might unintentionally signal an endorsement of automation, thereby threatening long-standing norms around authorship, creativity, and editorial authority. $P13_{\text{male, editor/journalist}}$ expressed this dilemma: “If we say it’s okay to use GenAI, would that encourage journalists to rely on it too much, instead of their own judgment? And how would our audience perceive this? Maybe they’d think, ‘Oh, this newspaper is just letting AI write the news now,’ or ‘Their journalism could be so poor that AI can step in.’ Some news audiences still carry a lot of negative impressions about GenAI.” This institutional hesitation creates a structural vacuum: GenAI becomes a tolerated but unregulated practice, operating in a gray zone between tacit acceptance and formal prohibition.
@@
Compounding this uncertainty, even newsrooms that recognize potential collaborative value often lack the expertise, confidence, or conceptual frameworks needed to operationalize GenAI at the teamwork level. As senior manager $P3_{\text{male, editor/manager}}$ put it bluntly: “We don’t even know how to try. What does GenAI collaboration actually mean?” $P3_{\text{male, editor/manager}}$ contrasted this ambiguity in collaboration with the current state of highly individualized GenAI use, where accountability remains clear. “Look, if everyone is using GenAI in a personal and unofficial way, and something goes wrong, we know who to hold responsible: it’s the individual journalist, right?” However, $P3_{\text{male, editor/manager}}$ raised concerns about what might happen if the newsroom were to formalize guidelines for how GenAI could be coordinated into team workflows. “Now imagine we create an official framework for collaboration between human journalists and GenAI. Let’s say we even publish a clear policy about how GenAI shall be incorporated into teamwork. Then if a problem arises, who’s accountable? The journalist might say, ‘I was just following the policy.’ So where does that leave us?”
These concerns illustrate a core tension facing many newsrooms. On the one hand, organizations hesitate to define GenAI collaboration for fear of enabling blame deflection, complicating accountability, or introducing institutional risk. On the other hand, leaving GenAI use uncoordinated preserves individual accountability but fragments practices, creating inconsistent, opaque, and difficult-to-scale workflows. As a result, many newsrooms remain trapped in a structural limbo—aware of GenAI’s potential to reshape journalistic collaboration, yet uncertain or unwilling to articulate what responsible, organization-level coordination would actually entail.
@@
6.1.2. Second, the technical infrastructure of most Chinese newsrooms poses significant challenges to building collaborative GenAI norms.
@@
Most GenAI tools were accessed individually via external platforms, such as Qwen Chat, Doubao, or DeepSeek. There are no built-in mechanisms in newsrooms to track when, where, or how GenAI has been used in the production of a specific story. As $P2_{\text{female, editor/manager}}$ explained, “People are mostly using GenAI on external platforms, which makes it hard for us to trace how they’re using it. We’re trying to build in-house GenAI systems with technical companies, but those are still in pilot stages. In most cases, journalists are still using external GenAI tools.”